OpenStudy (anonymous):

11/2 of what number is 6

OpenStudy (confluxepic):

\[\frac{ 11 }{ 2 }(x)=6\]

OpenStudy (confluxepic):

Multiply by 2/11 on both sides.

OpenStudy (confluxepic):

\[\frac{ 2 }{ 11 }\times\frac{ 11 }{2}(x)=6 \times \frac{ 2 }{ 11 }\]

OpenStudy (confluxepic):

\[x=\frac{ 12 }{ 11}\]